# Enhancement of Electric Drive in Silicon Quantum Dots with Electric Quadrupole Spin Resonance

### v2

@@ III Results and discussion
We first examine the results of the 5-electron regime, with the resulting energy levels shown in Fig. 4(a). Here, we are largely concerned with the $P_{x}$ and $P_{y}$ orbitals, given that we expect the valence electron to be sitting in the $P_{x}$ orbital based on electron-filling principles. We can obtain the qubit frequencies by calculating the energy difference between the ground and the first excited state, which we plot in Fig. 4(b). Here, the theoretical estimate is presented in red, contrasting with the experimental data in black. Finally, we can calculate the resulting Rabi frequency based on these qubit parameters, which is plotted in Fig. 4(c). The solid and dashed blue lines indicate respectively the Rabi frequency with and without EQSR contributions. It is important to highlight that this was plotted using the parameters fitted for the model that included EQSR, so it is trivial that the line representing only EDSR would not be a good fit. Our conclusions arrive from the fact that we tried as well to do a fitting procedure that does not include EQSR, but we could not reproduce the enhancement that follows the experimental data near the degeneracy point. This supports our theoretical hypothesis that EQSR is required to explain the Rabi frequency speed-ups.
We obtained similar results in the 13-electron regime, with the main difference being the orbitals on which the electric drive is acting, which are the $D$ orbitals in this case. We show the $D$ orbital energy levels, the qubit frequency fits, and the Rabi frequency fits in Figs. 4(d), (e), and (f) respectively. In this case, the qubit frequency is similarly defined as the difference between the ground and first excited energy of the $D$ orbital energy levels. In Fig. 4(f), we show that the enhancement in Rabi frequency can be mapped by the Hamiltonian with the inclusion of EQSR.
Therefore, in both the 5 electron and 13 electron cases, we demonstrate our ability to capture the behavior of the Rabi frequency speed-up. We stress that a quantitative correspondence between the experimental data and the theoretical fits is not the scope of this paper, and the main point is rather to show how the electrical quadrupole is necessary to explain the Rabi enhancement near the degeneracy point. The challenges in these fittings arise from several factors, including, but not limited to, insufficient experimental data points and having a multitude of parameters. We will discuss the technical aspects of the fitting protocol in greater detail in the Supplementary Material.
@@
Finally, we see from Eq. (9) that while Rabi frequency promoted by EDSR depends on $a_{0(1)}\propto\alpha+\beta$, the Rabi enhancement promoted by EQSR depends on $b_{0(1)}\propto\alpha^{2}-\beta^{2}$. This means that in order to improve Rabi frequencies next to the degeneracy point using intrinsic spin-orbit coupling, the latter relation between Rashba and Dresselhaus coefficients should be looked upon. If the difference between the square of the two terms are large enough, the addition of a micromagnet could be rendered unnecessary, which is desirable for scale-up architectures.
The experimental results in Ref. [26] show a clear correlation between the nonlinear bend of the qubit frequencies and the one order of magnitude increase in the Rabi frequencies of $P$ and $D$ orbitals. As we have shown here, both features can be explained within an effective single-particle theory if the quadrupole coupling is being taken into account. Our fits show the rising of Rabi frequency as dot shape becomes more circular, i.e. as we approach the anti-crossing point, therefore indicating clearly the role of EQSR in enhancing the Rabi frequency close to the degeneracy as a function of the ellipticity. In summary, we believe that the technique of quadrupole spin resonance opens up new pathways for coherent spin control with multi-electron dots and can become an ally in the quest for scalable quantum computation.

### v3

@@ III Results and discussion
We first examine the results of the 5-electron regime, with the resulting energy levels shown in Fig. 4(a). Here, we are largely concerned with the $P_{x}$ and $P_{y}$ orbitals, given that (for $\delta>1$) we expect the valence electron to be sitting in the $P_{y}$ orbital based on electron-filling principles (and in the $P_{x}$ orbital for $\delta<1$). We can obtain the qubit frequencies by calculating the energy difference between the ground and the first excited state, which we plot in Fig. 4(b). Here, the theoretical estimate is presented in red, contrasting with the experimental data in black. Finally, we can calculate the resulting Rabi frequency based on these qubit parameters, which is plotted in Fig. 4(c). The solid blue line indicate the Rabi frequency found accounting EDSR and EQSR, while the dashed blue line singles out the EDSR contribution. We also tried to do a fitting procedure that does not include EQSR, but we could not reproduce the enhancement that follows the experimental data near the degeneracy point. This supports our theoretical hypothesis that EQSR is required to explain the Rabi frequency speed-ups. Physically this is due to the fact that with EDSR alone the Rabi frequency would present a maximum when there is an anticrossing in Fig. 4(a). Nonetheless, we observe a maximum where the Stark shift (given by the slope in Fig. 4(b) is maximum, a feature only accounted by the EQSR coupling, as we can see in Fig. 4(c). Furthermore, the peak in the dashed lines of Figs 4(c) and (f) can be shown to be a higher order effect [20].
We obtained similar results in the 13-electron regime, with the main difference being that the electric drive is acting on the $D$ orbitals. We show the $D$ orbital energy levels, the qubit frequency fits, and the Rabi frequency fits in Figs. 4(d), (e), and (f) respectively. The qubit frequency is similarly defined as the difference between the ground and first excited energy of the $D$ orbital energy levels. In Fig. 4(f), we show that the enhancement in Rabi frequency can be mapped by the Hamiltonian with the inclusion of EQSR.
Therefore, in both the 5 electron and 13 electron cases, we demonstrate our ability to capture the behavior of the Rabi frequency speed-up. We stress that a quantitative correspondence between the experimental data and the theoretical fits is not the scope of this paper, but we show how the electrical quadrupole is necessary to explain the Rabi enhancement near the degeneracy point. The challenges in these fittings arise from several factors, including, but not limited to, insufficient experimental data points and having a multitude of parameters.
@@
Finally, we see from Eq. (9) that while Rabi frequency promoted by EDSR depends on $a_{0(1)}\propto\alpha+\beta$, the Rabi enhancement promoted by EQSR depends on $b_{0(1)}\propto\alpha^{2}-\beta^{2}$, which is key relation to consider in order to improve Rabi frequencies next to the degeneracy point using intrinsic spin-orbit coupling. If the difference between the square of the two terms are large enough, the addition of a micromagnet could become unnecessary, which is desirable for scale-up architectures. The EQSR theory not only explains behavior in previous experiments, but is also a formalism that accounts for phenomenon in the electrical control of quantum dots qubits.
The experimental results in Ref. [17] show a clear correlation between the nonlinear bend of the qubit frequencies and the one order of magnitude increase in the Rabi frequencies of $P$ and $D$ orbitals. Both features can be explained within an effective single-particle theory by taking quadrupole coupling into account. Our fits show the rising of Rabi frequency as dot shape becomes more circular, i.e. as we approach the anti-crossing point, therefore indicating clearly the role of EQSR in enhancing the Rabi frequency close to the degeneracy as a function of the ellipticity.
This work opens the perspective of analyzing how the EQSR combined with other coupling affects the driving mechanism, for example with valley-orbit coupling, which has recently been shown to perform EDSR[27]. Another perspective is to analyze the influence of electron-electron interaction on the EQSR mechanism. In summary, we believe that the technique of quadrupole spin resonance opens up new pathways for coherent spin control with multi-electron dots and can become an ally in the quest for scalable quantum computation.